In my math class we are doing Functions right now. The specific three we are using are Linear, Quadratic, and Exponential. Linear is y=mx+b, Quadratic is y=ax^2(+b) and Exponential is y=a*b^x. Each problem gives ordered pairs, or a table. I'll show you a table, and my answer, which I do not believe to be correct, since when I plug in some numbers from the table, doesnt not equal out.
EX:
X Y
-2 1.5
0 .5
2 -.5
4 -1.5
y=-x +.5
Now, that is how my math teacher taught us to solve it; find the slope, y intercept when x=0, etc, and when you plug numbers in from the table, my answer does not work. Obviously, since you are very intelligent, you will be able to see my flaw. Thank you.

Answer by venugopalramana(3286)

IF YOU KNOW EQNS...STRAIGHT FORWARD APPROACH IS SOLVE FOR THE UNKNOWNS ASSUMING AN EQN.HERE YOU ARE TRYING LINEAR..OK...
FOR -2,1.5......WE GET
1.5=-2A+B.....FOR THE NEXT PAIR OF 0,0.5
0.5=2*0+B...OR...B=0.5
SO 1.5=-2A+0.5....OR......-2A=1.5-0.5=1....OR...A=-1/2=-0.5
SO EQN.IS
Y=-0.5X+0.5...NOW TEST THE OTHER PAIRS TO SEE IF THIS IS CORRECT....
-0.5=-0.5*2+0.5=-1+0.5=-0.5...OK
-1.5=-0.5*4+0.5=-2+0.5=-1.5...OK
SO THE EQN.IS LINEAR AND IT IS
Y=-0.5X+0.5